# Exploring Context-aware and LLM-driven Locomotion for Immersive Virtual Reality

### v1

@@ 3.3.2 IPQ Scores
The IPQ results, shown in Figure 8, present scores across four dimensions of presence: general presence, spatial presence, realism, and involvement, for each locomotion condition. Scores range from 0 to 6. The LLM-driven locomotion method received the highest ratings in both general presence (, ) and spatial presence (, ), indicating that participants experienced a stronger sense of immersion in the virtual environment. In the realism dimension, the voice-based steering condition scored the highest (, ), suggesting that participants perceived the environment as more realistic when using voice command-based steering. For involvement, teleportation resulted in the highest score (, ), while the LLM-driven locomotion (, ) and voice-based steering methods (, ) showed similar values. We did not find statistically significant differences across the conditions ().
@@ 3.3.3 NASA-TLX Questionnaire
3.3.4 SUS Questionnaire
3.4 Model Evaluation and Feature Importances
We conducted a comprehensive analysis to evaluate the performance of various machine learning models and identify key eye-tracking features changing with the locomotion method. The baseline accuracy obtained by predicting the majority class fix command-based voice steering was . In the test set, the class distributions are 55 samples () for LLM-driven locomotion, 56 samples () for fix-command-based voice steering, and 19 samples () for teleportation. We summarize the classification accuracies of each model in Table 3. Among these, LightGBM achieved the highest performance, with an accuracy of , followed by Gradient Boosting and Random Forest.
To interpret the effect of locomotion techniques on eye movement characteristics, we used SHAP to evaluate the contribution of individual features within the LightGBM classifier. The SHAP summary plot is shown in Figure 11, illustrating the average impact of each feature on the model’s output, aggregated by classes teleportation, voice-based steering through fixed commands, and LLM-driven locomotion. According to the SHAP analysis, total fixation duration, minimum pupil size, and mean saccade duration emerged as the most important features for distinguishing between locomotion techniques. These features are commonly linked to cognitive load, as well as visual attention, engagement, and visual search behavior, suggesting that user interaction patterns varied across conditions. Additionally, total saccade duration and maximum blink duration contributed substantially to the model’s predictions, indicating further differences in attentional strategies. Other influential features included the standard deviation of saccade duration, minimum saccade peak velocity, and blink count. Overall, the results indicate that fixation behavior, saccade dynamics, blink activity, and pupil-based measures collectively shape the gaze responses associated with each locomotion technique, reflecting distinct patterns of cognitive processing and user engagement in VR.
@@ 4.1 Performance and User Experience
Teleportation, while the fastest method, relies on hand controllers, making it unsuitable for hands-free scenarios. Among the hands-free methods, voice-based steering initially enabled faster navigation to the first target, but participants quickly adapted to the LLM-driven method, resulting in more efficient performance in the second phase and faster overall completion. This learning effect suggests that users quickly adapted to the natural language system and were able to make effective use of its flexibility, despite being completely unfamiliar with this form of locomotion.
@@
The LLM-driven method, which enables instant transport to target locations, reduced nausea and vestibular discomfort compared to continuous movement in voice-based steering. This aligns with prior work showing that continuous motion is more likely to induce cybersickness [8]. Although all participants reached the same targets, those using teleportation completed the task much faster, resulting in significantly less exposure to the virtual environment—nearly four times shorter than other conditions. This reduced duration likely contributed to the lower overall cybersickness scores in the teleportation condition. Interestingly, the pattern reversed for oculomotor discomfort, where voice-based steering yielded slightly lower scores. This may indicate that its steady, continuous movement places less strain on the visual system than the abrupt position changes used in the other techniques.
Teleportation required the least cognitive effort, while the LLM-driven method showed slightly higher cognitive demand than voice-based steering. A similar trend was observed in the pupil diameter measurements. The increase in the LLM-driven condition may reflect the mental effort required to plan contextually relevant and efficient commands, which is not expected in teleportation or in command-based methods that rely on a simple and repeatable set of inputs. However, it may also promote deeper cognitive engagement and encourage more active exploration of the immersive virtual environment, potentially contributing to the heightened sense of presence reported in the LLM-driven locomotion condition. Additionally, although free-form commands may initially impose a higher cognitive load, this demand is likely to decrease with experience, which can be further investigated through longitudinal studies. Despite the slightly higher cognitive load associated with contextual commands, participants using LLM-driven locomotion reported the highest levels of general and spatial presence. Actively engaging with the environment to provide context-aware instructions likely strengthened their sense of being “present” compared to voice-based steering. In contrast, voice-based steering received the highest realism scores, possibly because its continuous walking pattern more closely mimics real-world locomotion.
@@ 4.2 Cognitive and Visual Attention
Voice-based steering resulted in significantly higher fixation rates, indicating increased visual scanning and attentional demands during continuous navigation. This result may be attributed to the nature of constant movement, where users need to actively monitor their direction and rapidly search the environment, often resulting in shorter and more frequent fixations. Similarly, mean fixation duration serves as an indicator of stable visual attention and cognitive engagement. Teleportation resulted in the longest fixations, followed by the LLM-driven method, while the voice-based steering condition showed the shortest durations. These eye movement patterns, particularly less frequent but longer fixations in the teleportation and LLM-driven conditions, suggest that users were able to plan their actions more deliberately and were generally aware of their intended destinations.
@@
Classification models demonstrate the feasibility of using eye-tracking features to infer user interaction mode in real time, as these features varied across locomotion techniques. SHAP analysis indicated that fixation duration, saccade dynamics, and pupil diameter were key indicators for distinguishing user behavior. These distinct eye movement patterns, which reflect variations in visual attention and cognitive engagement, can serve as objective metrics for adaptive VR systems.
@@ 4.3 Accessibility and Practical Considerations
The LLM-driven locomotion method provides a highly accessible alternative for users with disabilities or temporary impairments, matching teleportation and fixed-command methods in usability and comfort. With its inherent multilingual support, users can interact naturally in their own languages without relying on language-specific phrasing. This eliminates the need for environment-specific adaptations or manual configuration, making the system highly adaptable and inclusive. Such flexibility is especially valuable in settings like medical training, education, and remote collaboration, where accessibility and ease of use are critical as VR adoption grows.
However, one challenge in implementing LLM-driven locomotion in real-world scenarios is the inherently dynamic and potentially unpredictable nature of LLM outputs, despite their strong performance across a wide range of conditions. Since they are not entirely controllable, ensuring consistent and predictable behavior may require additional error-handling mechanisms. These may include constraining the range of movement, validating target coordinates, or keeping the user in place when the model output is ambiguous or refers to non-navigable areas. Another challenge is finding the right balance between accuracy and performance in LLM and speech-to-text models. While reducing latency by using smaller models can improve responsiveness, it is crucial to maintain strong overall performance.
@@ 5 Conclusion
In this work, we proposed an LLM-driven, hands-free locomotion technique for VR that uses natural language instructions to enable instant teleportation, enhancing accessibility and addressing the motion sickness often associated with continuous voice-based steering. To assess the effectiveness of this method, we conducted a comparative evaluation of three locomotion techniques: teleportation, voice-based steering through fixed commands, and our LLM-driven approach. We used eye-tracking analysis to gain insight into users’ attention, engagement, and cognitive processes and complemented this analysis with standardized questionnaires, including the SUS, IPQ, CSQ-VR, and NASA-TLX, to assess usability, presence, cybersickness, and cognitive workload, respectively. While teleportation remains the most efficient locomotion method, it is often not suitable for hands-free scenarios. The proposed LLM-driven approach offers a promising alternative by combining natural language interaction with contextual awareness and instant movements while also inherently supporting multiple languages as an advantage. Although it requires slightly more cognitive effort, it enhances immersion and comfort, particularly in terms of reducing cybersickness. Overall, our results demonstrate that LLM-driven locomotion is a promising alternative for intuitive, accessible, and immersive hands-free navigation in VR environments. Future research could investigate personalization techniques, such as adapting LLM behavior based on user preferences or prior interactions, to enhance predictability and usability over time.

### v2

@@ 3.3.2 IPQ Scores
The IPQ results, shown in Figure 8, present scores across four dimensions of presence: general presence, spatial presence, realism, and involvement, for each locomotion condition. Scores range from 0 to 6. The LLM-driven locomotion method received the highest ratings in both general presence (, ) and spatial presence (, ), indicating that participants experienced a stronger sense of immersion in the virtual environment. In the realism dimension, voice-based steering scored highest (, ), suggesting participants perceived the environment as more realistic with this method. For involvement, teleportation resulted in the highest score (, ), while the LLM-driven locomotion (, ) and voice-based steering (, ) showed similar values. However, we found no statistically significant differences across the conditions ().
@@ 3.3.3 NASA-TLX Questionnaire
3.3.4 SUS Questionnaire
3.4 Model Evaluation and Feature Importances
We conducted an exploratory analysis to evaluate the performance of various machine learning models and identify key eye-tracking features changing with the locomotion method. The baseline accuracy obtained by predicting the majority class fix command-based voice steering was . In the test set, the class distributions are 55 samples () for LLM-driven locomotion, 56 samples () for fix-command-based voice steering, and 19 samples () for teleportation. We summarize the classification accuracies of each model in Table 3. Among these, LightGBM achieved the highest performance, with an accuracy of , followed by Gradient Boosting and Random Forest.
@@ 3.4 Model Evaluation and Feature Importances
To interpret the effect of locomotion techniques on eye movement characteristics, we used SHAP to evaluate the contribution of individual features within the LightGBM classifier. The SHAP summary plot is shown in Figure 11, illustrating the average impact of each feature on the model’s output, aggregated by classes teleportation, voice-based steering through fixed commands, and LLM-driven locomotion. According to the SHAP analysis, total fixation duration, minimum pupil size, and mean saccade duration emerged as the most important features for distinguishing between locomotion techniques. These features are commonly linked to cognitive load, as well as visual attention, engagement, and visual search behavior, suggesting that user interaction patterns varied across conditions. Additionally, total saccade duration and maximum blink duration substantially contributed to the model’s predictions, indicating further differences in attentional strategies. Other influential features included saccade duration variability, minimum saccade peak velocity, and blink count. Overall, the results indicate that fixation behavior, saccade dynamics, blink activity, and pupil-based measures collectively shape the gaze responses associated with each locomotion technique, reflecting distinct patterns of cognitive processing and user engagement in VR.
@@ 4.1 Performance and User Experience
Teleportation, while the fastest method, relies on hand controllers, making it unsuitable for hands-free scenarios. Among the hands-free methods, voice-based steering initially enabled faster navigation to the first target, but participants quickly adapted to the LLM-driven method, which led to faster overall completion as shown in Section 3.1. This learning effect suggests users quickly adapted to the natural language system and effectively used its flexibility, despite being completely unfamiliar with this form of locomotion.
@@
The LLM-driven method, which enables instant transport to target locations, caused lower nausea and vestibular discomfort scores than voice-based steering, though the differences were not statistically significant. This aligns with prior work showing that continuous motion is more likely to induce cybersickness [8]. Although all participants reached the same targets, those using teleportation completed the task much faster, resulting in significantly less exposure to the virtual environment—nearly four times shorter than other conditions. This reduced duration likely contributed to the lower overall cybersickness scores in the teleportation condition. Interestingly, the pattern reversed for oculomotor discomfort, with voice-based steering scoring slightly lower, suggesting that continuous movement may strain the visual system less than abrupt position changes.
Teleportation required the least cognitive effort, while the LLM-driven method showed slightly higher demand than voice-based steering, although these differences were not statistically significant. A similar trend was observed in the pupil diameter measurements. The increase in the LLM-driven condition may reflect the mental effort required to plan contextually relevant and efficient commands, which is not expected in teleportation or in command-based methods that rely on a simple and repeatable set of inputs. However, it may also promote deeper cognitive engagement and encourage more active exploration of the immersive virtual environment, potentially contributing to the heightened sense of presence reported in the LLM-driven locomotion condition. Additionally, although free-form commands may initially impose a higher cognitive load, this demand is likely to decrease with experience, which can be further investigated through longitudinal studies. Despite the slightly higher cognitive load associated with contextual commands, participants using LLM-driven locomotion reported the highest levels of general and spatial presence. Actively engaging with the environment to provide context-aware instructions likely strengthened their sense of being “present” compared to voice-based steering. In contrast, voice-based steering received the highest realism scores, possibly because its continuous walking pattern more closely mimics real-world locomotion.
@@ 4.2 Cognitive and Visual Attention
Voice-based steering resulted in significantly higher fixation rates, indicating increased visual scanning and attentional demands during continuous navigation. This result may be attributed to the nature of constant movement, where users need to actively monitor their direction and rapidly search the environment, often resulting in shorter and more frequent fixations. Similarly, mean fixation duration serves as an indicator of stable visual attention and cognitive engagement. Teleportation resulted in the longest fixations, followed by the LLM-driven method, while the voice-based steering condition showed the shortest durations. These eye movement patterns, particularly less frequent but longer fixations in the teleportation and LLM-driven conditions, suggest that users planned their actions more deliberately and were generally aware of their intended destinations.
@@
Classification models demonstrate the feasibility of using eye-tracking features to infer user interaction mode in real time, as these features varied across locomotion techniques. SHAP analysis indicated that fixation duration, saccade dynamics, and pupil diameter were key indicators for distinguishing user behavior. Such distinct eye movement patterns, reflecting visual attention and cognitive engagement, can serve as objective metrics for adaptive VR systems.
@@ 4.3 Accessibility and Practical Considerations
The LLM-driven locomotion method provides a highly accessible alternative for users with disabilities or temporary impairments, as it does not require additional hardware such as handheld controllers, and matches teleportation and fixed-command methods in usability and comfort. With its inherent multilingual support, users can interact naturally in their own languages without relying on language-specific phrasing. This advantage eliminates the need for environment-specific adaptations or manual configuration, making the method highly adaptable and inclusive. It offers a plug-and-play approach, as it can be easily generalized to virtual environments. Such flexibility is especially valuable in settings like medical training, education, and remote collaboration, where accessibility and ease of use are critical as VR adoption grows.
At the same time, several practical challenges should be considered. Handling variability in user prompts requires careful prompt design and contextual grounding, particularly for real-time proper responses. In our implementation, the system prompt incorporated essential scene context (see Figure Exploring Context-aware and LLM-driven Locomotion for Immersive Virtual Reality), enabling the LLM to interpret intent accurately and generate only the target coordinates. Outputs are short, and responses average 1.5 seconds without disrupting interaction or causing discomfort, as the movement is not continuous. Another challenge is the inherently dynamic and sometimes unpredictable nature of LLM outputs. Ensuring consistent behavior in real-world scenarios requires error-handling mechanisms, such as constraining movement ranges, validating coordinates, or ignoring ambiguous outputs. Although smaller LLMs could be used, prompt design is more critical in such cases, as larger models are generally more reliable in capturing context and producing properly formatted outputs. Additionally, our approach represents the scene as text, enabling real-time processing but potentially overlooking spatial layouts. As scene complexity increases, larger prompts may introduce latency and reduce robustness. Nevertheless, LLMs interpret such inputs accurately with well-designed prompts.
@@ 5 Conclusion
In this work, we proposed an LLM-driven, hands-free locomotion technique for VR that uses natural language instructions, enhancing accessibility and addressing the motion sickness often associated with continuous voice-based steering. To assess its effectiveness, we compared three locomotion techniques: teleportation, voice-based steering through fixed commands, and our LLM-driven approach. We used eye-tracking analysis to gain insight into users’ attention, engagement, and cognitive processes and complemented this analysis with standardized questionnaires, including the SUS, IPQ, CSQ-VR, and NASA-TLX, to assess usability, presence, cybersickness, and cognitive workload, respectively. While teleportation remains the most efficient locomotion method, it is often not suitable for hands-free scenarios. The proposed LLM-driven approach offers a promising alternative by combining natural language interaction with contextual awareness and instant movements while also inherently supporting multiple languages as an advantage. Although it requires slightly more cognitive effort, it enhances immersion and comfort, particularly in terms of reducing cybersickness. Overall, our results demonstrate that LLM-driven locomotion is a promising alternative for intuitive, accessible, and immersive hands-free navigation in VR environments. Future research should investigate personalization techniques, such as adapting LLM behavior based on user preferences or prior interactions, to enhance predictability and usability over time.